Morrisey & Brown, Ltd., of Sydney is a merchandising company that is the sole distributor of a product that is increasing inRequired: 1. Identify each of the companys expenses (including cost of goods sold) as either variable, fixed, or mixed. Expe2. Using the high-low method, separate each mixed expense into variable and fixed elements. State the cost formula for each m3. Redo the companys income statement at the 5,000-unit level of activity using the contribution format. Morrisey & Brown, L
Answer #1
1) Expenses Classification
cost of goods sold Variable
Advertising expense fixed
shipping expense mixed
Salaries and commission mixed
insurance expense fixed
Depreciation expense fixed
2) High low method
shipping expense
units cost
high 5,000 38,000
low 4,000 34,000
difference 1,000 4,000
variable expense per unit 4000/1000 = $4
fixed expense =38000-5000*4 =18000
Salaries and commission
units cost
high 5,000 90,000
low 4,000 78,000
difference 1,000 12,000
Variable expense per unit = 12000/1000=$12.00
fixed expense = 90000-5000*12 =30000
cost formula
Variable cost Fixed cost                    Formula                                                       
shipping expense 4 per unit 18,000 Y= 18,000 + 4 per unit
Salaries & commission 12 per unit 30,000 Y= 30,000 + 12 per unit
3) Sales in units 5000
Sales 500,000
Variable expense
Cost of goods sold 300,000
Shipping expense 20,000
Salaries and commission 60000
380,000
Contribution margin 120,000
Fixed expense
Advertising expense 21,000
Shipping expense 18,000
Salaries & commission 30,000
insurance expense 6,000
Depreciation expense 15,000
90,000
Net income 30,000